OpenStudy (haleyelizabeth2017):

Jana hit a golf ball with an initial velocity of 102 feet per second at an angle of 65 degrees with the horizontal. After 2 seconds, how far has the ball traveled horizontally?

OpenStudy (isaidavila):

Decompose the velocity vector vertically and horizontally. The horizontal vector is Vcos© so V1=Vcos65 V1=102*cos65 Distance is velocity times time D=V1*t D=Vcos65*2 D=102cos65*2 Just solve

OpenStudy (haleyelizabeth2017):

I don't quite understand..